OpenStudy (anonymous):

Suppose you are traveling straight from Los Angeles, California to Albuquerque, New Mexico, stopping overnight in Flagstaff, Arizona. ... If the distance from L.A. to Flagstaff is 120 miles greater than the distance from Flagstaff to Albuquerque, how far is L.A. from Flagstaff, and how far is Flagstaff from Albuquerque?

OpenStudy (yanasidlinskiy):

Let x be the distance between Flagstaff and Albuquerque. Solve the following for x: (120 + x) + x = 810 x = 345 miles LA to Flagstaff is (120 + x) = (120+345) = 465 miles.

OpenStudy (anonymous):

thats how i did it. just wanted to make sure i was right thanks.

OpenStudy (yanasidlinskiy):

No problem!!:) Glad I could help!:)